OpenStudy (butterfield1215):

4+3x=y what is the slope with fan and Medal

HanAkoSolo (jamierox4ev3r):

Hi @Butterfield1215 , the easiest way to find the slope is to solve for y The goal is to get this equation in slope intercept form, which is \(y=mx+b\)

OpenStudy (butterfield1215):

ok

HanAkoSolo (jamierox4ev3r):

In this equation, however, we already solved for y. we just need to rearrange the equation. y=3x+4

OpenStudy (butterfield1215):

so its 3/1

OpenStudy (♪chibiterasu):

Yeah. You don't need to make it a fraction.

OpenStudy (butterfield1215):

oh ok

OpenStudy (♪chibiterasu):

\[\frac {3}{1} = 3\]

HanAkoSolo (jamierox4ev3r):

But in the equation \(y=mx+b\) \(\ast\)m is the slope \(\ast\)b is the y intercept. so the value next to x is the slope. and you don't need to make it a fraction :)

HanAkoSolo (jamierox4ev3r):

Does that make sense to you? anytime, given and equation in \(y=mx+b\) format, the \(m\) is ALWAYS the slope :)

OpenStudy (butterfield1215):

I understand what your saying